OpenStudy (anonymous):

Solve: 5(x − 1) = 3(x + 2) − 7

OpenStudy (lncognlto):

5(x - 1) = 3(x + 2) - 7 5x - 5 = 3x + 6 - 7 2x = 11 - 7 2x = 4 x = 2

OpenStudy (lncognlto):

They aren't? 5(x-1) = 5x - 5. 3(x + 2) = 3x + 6. Moving all the x values to the left and all the numerical values to the right gives 5x - 3x = 6 + 5 - 7. Thus 2x = 11 - 7 therefore 2x = 4 Therefore x = 2.

OpenStudy (uri):

5x-5=3x+6-7 5x-3x=6-7+5 2x = 11 - 7 2x=4 x=4/2 x=2 Sorry bloody typo :P

OpenStudy (lncognlto):

Hahaha - Yes, I know those ;)